If tantheta +cottheta =2Then tan^7theta +cot^7theta =?

Grade:10

1 Answers

Arun
6 years ago
Dear student
Let tan theta = x
Then
x + 1/x = 2
This is only possible when x = 1
Hence
Tan theta = 1 and also cot theta = 1
 
Hence
tan^7 theta + cot^7 theta
= 1^7 + 1^7
= 1 +1
= 2
 
Regards
Arun (askIITians forum expert)